OpenStudy (anonymous):

SOMEONE HELP ASAP!!!!! WILL FAN AND MEDAL!!!!!!!! 1. 4 - t = 3(t-1) - 5 A. 3 B. 6 C. no solution D. identity 2. 8x - 2(x + 1)= 2(3x - 1) A. 0 B. 2 C. no solution D. identity 3. 3(c - 2) = 2(c - 6) A. 2 B. -6 C. no solution D. identity 4. 0.5(m + 4) = 3(m - 1) A. -2 B. 2 C. no solution D. identity

OpenStudy (shinalcantara):

4-t=3(t-1)-5 4-t=3t-3-5 -4t=-12 t=3 now try to solve for the others applying the same concept

OpenStudy (anonymous):

im confused with the others, could you help me work them out?